# Polymer Collapse & Liquid-Liquid Phase-Separation are Coupled in a Generalized Prewetting Transition

### v1

@@ Abstract
The three-dimensional organization of chromatin is thought to play an important role in controlling gene expression. Specificity in expression is achieved through the interaction of transcription factors and other nuclear proteins with particular sequences of DNA. At unphysiological concentrations many of these nuclear proteins can phase-separate in the absence of DNA, and it has been hypothesized that, in vivo, the thermodynamic forces driving these phases help determine chromosomal organization. However it is unclear how DNA, itself a long polymer subject to configurational transitions, interacts with three-dimensional protein phases. Here we show that a long compressible polymer can be coupled to interacting protein mixtures, leading to a generalized prewetting transition where polymer collapse is coincident with a locally stabilized liquid droplet. We use lattice Monte-Carlo simulations and a mean-field theory to show that these phases can be stable even in regimes where both polymer collapse and coexisting liquid phases are unstable in isolation, and that these new transitions can be either abrupt or continuous. For polymers with internal linear structure we further show that changes in the concentration of bulk components can lead to changes in three-dimensional polymer structure. In the nucleus there are many distinct proteins that interact with many different regions of chromatin, potentially giving rise to many different Prewet phases. The simple systems we consider here highlight chromatin’s role as a lower-dimensional surface whose interactions with proteins are required for these novel phases.
Introduction
@@
Long biopolymers, DNA & RNA, encode genetic information in their sequence. The three dimensional structure and arrangement of these polymers plays an important role in cellular processes. In the nucleus of eukaryotes, the chromosome displays three dimensional structure at many distinct scales, from TADs to euchromatin/heterochromatin [1], and in some cases the co-localization of enhancers, promoters, and otherwise distal genetic elements has been shown to be a key step in initiating transcription [2, 3, 4]. While some of this structure is driven by enzymatic activity from loop extrusion factors (LEFS), helicases, isomerases, and polymerases  [5, 6], some is likely to be thermodynamic in origin, driven by energetic interactions between long polymers and proteins, small RNAs and other macromolecules [7, 8].
Many nuclear components including transcription factors and heterochromatin-associated proteins have a tendency to phase separate into coexisting liquid phases when isolated at high concentrations, even in the absence of DNA [9, 7, 10]. This phase separation is likely driven by large numbers of weak but specific interactions between these components, which are sufficiently strong to overcome the entropic cost of phase separating, but weak enough to allow for mobility of individual components. Phase separation of protein components is a driving force behind a range of functional structures in the cytoplasm, and has been argued to play a similar role in organizing chromatin [11, 12, 2, 13].
@@
But there are other thermodynamic drivers of spatial organization available to long polymers like chromosomes and RNA. In particular, long polymers can undergo thermodynamic transitions of their own. Collapse transitions separate a regime where typical configurations are extended from where polymers are condensed and space filling [14]. These transitions can be accessed by solvent quality; in good solvents polymers typically have a weak contact repulsion and exist in an Extended configurational phase. In poor solvents, polymers prefer to minimize their solvent interaction by maximizing self-contacts, leading to Collapsed phases. In the nucleus more complex interactions could in principle drive more interesting configurational phases.
@@
We recently argued that a range of structures at the plasma membrane are Prewet: surface phases formed by a combination of membrane mediated forces and interactions between cytoplasmic proteins [15]. In classical prewetting, a thin film resembling a three dimensional phase can adhere to a surface outside of a thermodynamic regime where that bulk phase is stable [16, 17]. But in the biological case, the solid surface is replaced by the fluid plasma membrane, itself found near a 2D liquid-liquid demixing critical point [18]. We showed that in this system the prewetting transition of the bulk merges with the de-mixing transition at the surface, allowing for phase transitions that would not occur if the membrane or the bulk were isolated from each other. This Prewet phase can also be viewed from the perspective of the membrane with bulk proteins shifting the location of the two dimensional liquid demixing transition.
@@
Here we explore a model in which bulk proteins prone to phase separation play an analogous role modulating the properties of a polymer collapse transition. In this generalized prewetting transition we consider a long polymer in a good solvent that interacts with bulk components. As we will show, this system has a phase in which bulk components are at near condensed density, and in which the polymer is Collapsed, at parameters in which the polymer would remain Extended in the absence of bulk, and the bulk would remain well-mixed in the absence of a polymer to scaffold it. We investigate the phase diagram of this generalized prewetting transition with both a lattice model and a mean field theory, and explore its implications for genomic structure.
Results
@@
Model — We model a single self-avoiding polymer on a $3$-D cubic lattice of linear dimension $L=64$. A spin variable at lattice site $i$, $s^{p}_{i}=1$ for sites which the polymers passes through and is 00 otherwise. The energy of configurations of this system is given by the Hamiltonian $H_{\text{poly}}$:
@@
Here both sums are over lattice sites, and $<i,j>$ represents a sum over nearest neighbors in space. $\mu_{p}$ is a monomer chemical potential which multiplies $N_{\text{p}}=\sum_{i}s_{i}^{\text{p}}$, the total number of monomers in the polymer. $v_{\text{p}}$ is a monomer-monomer interaction energy, and the subtraction in the first term avoids counting contacts between sequential monomers. All simulations take place on periodic lattices, and are constrained to polymer configurations with initial position $(L/2,L/2,0)$ and final position $(L/2,L/2,L)$. We also enforce that the polymer have length shorter than $N_{\text{max}}$, typically $1500$ monomer units, so that the polymer obeys $L\leq N_{\text{p}}<N_{\text{max}}$. In most of our simulations $N_{\text{p}}\approx N_{\text{max}}$, so that we effectively sample a fixed length distribution, but we allow small fluctuations for numerical convenience.
We couple this single long polymer (yellow in Figure 1A) to a solution of two types of shorter polymers which we term bulk molecules (red and blue in Figure 1A) and which are held at fixed chemical potential $\mu_{\text{b}}$ and fixed length $N_{\text{b}}$. Following our previous work [15], we model the bulk as a mixture of two types of molecules prone to condensing together. Molecules of the same type are self-avoiding while molecules of different types may occupy the same site, and have an on-site attractive interaction with energy $J_{\text{bulk}}$. Their occupancies at site $i$ are defined analogously to the long polymer by two additional spin variables $s^{1}_{i}$ and $s^{2}_{i}$. All molecules interact with a weak nearest-neighbor energy $J_{\text{nn}}$. The bulk Hamiltonian thus reads:
$\frac{H_{\text{bulk}}}{k_{\text{B}}T}=$ $-\frac{\mu_{\text{b}}}{N_{\text{b}}}\sum_{i}\left(s^{1}_{i}+s^{2}_{i}\right)+J_{\text{bulk}}\sum_{i}s^{1}_{i}s^{2}_{i}$
$+J_{\text{nn}}\sum_{<i,j>}\left(s^{1}_{i}s^{1}_{j}+s^{1}_{i}s^{2}_{j}+s^{2}_{i}s^{2}_{j}\right)$ (2)
The bulk system corresponds to proteins such as Transcription Factors that interact with each other and with chromosomes. $J_{\text{bulk}}$ represents the interaction between these proteins, modulated by post-translational modifications such as phosphorylation. And a chemical potential $\mu_{\text{b}}$ sets the concentration of these proteins.
Figure 2: The bulk system shifts the Polymer Collapse phase boundary — (A-D) Configurations of coupled phases: Outside of the bulk coexistence region there are two phases that correspond to phases of an isolated polymer: (A) Extended and (B) Collapsed. The Collapsed phase coincides with condensation, or Prewetting, of the bulk fluid. The Extended phase has corresponding Dilute bulk phase. Within the bulk coexistence region the polymer may prefer to enter the droplet, confining itself (Wet, C) or avoiding it (Dry, D). (E) Schematic bulk phase diagram: Generalized Prewet phases occur outside of the bulk coexistence region and terminate in a critical point above the bulk critical temperature. Colored points correspond to above configurations. (F) Schematic phase diagram of monomer-monomer interactions $v_{\text{p}}$ and bulk chemical potential $\mu_{\text{b}}$. The bulk modifies the phase boundary of the polymer, allowing it to collapse in regions where it otherwise would not; this Collapse transition coincides with condensation of the bulk on the surface of the polymer.
@@
$J_{\text{int}}$ can be interpreted as the affinity of a protein for a specific DNA sequence or histone modification. With these interactions, the red bulk molecule could correspond to a transcription factor and the blue to a co-activator such as Mediator. We illustrate this model in Figure 1A.
@@
We sample equilibrium configurations of this system using a Monte-Carlo procedure, fully explained in the Methods. Briefly, we generate a long polymer through a set of three elementary moves corresponding to the addition, deletion, and ‘kink’ of a given bond. Bulk molecules are updated with reptation and kink moves that conserve length, and we exchange these particles with a reservoir. We accept moves with Metropolis probability chosen to obey detailed balance given the above Hamiltonian.
@@
Thermodynamics of an Isolated Polymer — In the absence of the bulk system, the long polymer can access several thermodynamic phases. These phases correspond to distributions over the spatial configurations of the long polymer, $\left\{\vec{r}_{i}\right\}$, where $\vec{r}_{i}$ is the position of the $i^{\text{th}}$ monomer. For idealized polymers in regions where boundary conditions can be neglected, these phases can be characterized by scaling relationships between $N_{\text{p}}$ and the radius of gyration $R_{\text{g}}$, defined by $R_{\text{g}}^{2}=1/N_{\text{p}}\sum_{i}^{N_{\text{p}}}\left(\vec{r}_{i}-\vec{r}_{\text{cm}}\right)^{2}$ with $\vec{r}_{cm}$ the center of mass of the chain. The relation $R_{g}\sim N_{\text{p}}^{\nu}$ defines the exponent $\nu$ which characterizes different phases. We show sample configurations of these phases in Figure 1B
@@
When $\mu_{p}$ is negative and $v_{\text{p}}$ is positive or small and negative, the polymer is in the Short phase. Here $N_{\text{p}}$ is small, $\approx L$, and the scaling exponent $\nu\approx 1$.; the polymer is a fluctuating chain of length $N$ whose end-to-end distance $R_{\text{g}}\sim N_{\text{p}}$. In our simulations we typically set $\mu_{\text{p}}=0$, and are not concerned with this Short phase.
@@
When $v_{\text{p}}$ is large and attractive, the polymer is in the Collapsed phase. Here $N_{\text{p}}$ fluctuates about $N_{\text{max}}$, and the scaling exponent $\nu$ approaches the space-filling limit of $\frac{1}{3}$ in $3D$. We show a typical Collapsed phase in  1B.
@@
When $v_{\text{p}}$ is large and repulsive, but $\mu_{p}$ is positive, the polymer is in the Extended phase. Here the length fluctuates about $N_{\text{max}}$, but the polymer configuration is Extended and the average density of the polymer approaches 00 as $N_{\text{max}}$ approaches infinity. The Extended and Collapsed phase are separated by a line of continuous transitions at $v_{\text{p}}=0$. We show a typical Extended configuration in 1B.
@@
Bulk Molecules Phase Separate and Modify the Polymer Phase Diagram — Our bulk system has phases and phase-transitions independent of its interactions with the long polymer. The bulk phases are dilute and dense corresponding to gas and liquid-like states. At certain chemical potentials, or densities, these bulk phases will coexist and the system phase-separates, drawn as the yellow region in Figure 2E, with typical configurations in each phase shown in Figure  1C.
@@
Within the bulk coexistence region the polymer can prefer to localize to the dense phase or the dilute phase. This transition occurs when the energy of polymer-bulk interactions surpasses the entropic and energetic penalties of confining the long polymer to the bulk dense phase. This transition is analogous to the Wetting transition: when $J_{\text{int}}$ is strong the polymer prefers to enter the droplet to maximize its interactions with the bulk dense phase. We show sample configurations from Dry and Wet phases in Figure 2C,D. The polymer configurations in the Wet phase depend on $J_{\text{int}}$ and $J_{\text{nn}}$, which together determine solvent quality of the dense phase. A polymer that prefers to be Extended in an infinite bulk dense phase can be confined by a finite-size droplet separate from the collapse transition.
@@
Outside of the bulk coexistence region we find another phase where the polymer is Collapsed and also enriched with a dense-like phase of bulk molecules, we show a typical configuration in 2B. This is reminiscent of prewetting where a thin bulk domain forms which is adhered to a particular surface phase [15, 17]; we call phases where polymer collapse coincides with polymer-only bulk de-mixing Generalized Prewet - Collapsed. Consistent with this view, the polymer phase-diagram effectively shifts and the collapse transition occurs at repulsive $v_{\text{p}}$. Polymer collapse and bulk demixing couple because this minimizes the volume of an otherwise unfavorable bulk phase, while maintaining a high density of bulk-bulk and bulk-polymer interactions. Increasing $\mu_{\text{b}}$, we transition discontinuously between the three phases shown in Figure  2A-C, which we draw schematically in the phase diagrams sketched in Figure 2E,F.
@@
Here we develop a mean-field theory that couples a long, collapsible polymer to a bulk fluid with a propensity to phase-separate. We begin by briefly re-deriving the phase-diagram of an isolated polymer chain, then explore the thermodynamics of the coupled system.
@@
The first two terms labeled ‘Non-Interacting polymer’ are entropic contributions to the free-energy from gaussian statistics, and can be derived by replacing the radius of gyration with the end-to-end distance, with $a$ the linear dimension of a single monomer. A free-energy written in terms of $R_{\text{g}}$ has additional entropic costs due to confinement – but these, along with higher order interactions, are never relevant so we neglect them here (see Supplement). The final two terms capture mean-field interactions between monomers [19]. The parameter $v_{\text{p}}$ is the difference between monomer-monomer and monomer-solvent interactions. The point $v_{\text{p}}=0$ separates an Extended polymer phase, where the monomer density $\rho=Na^{3}/R_{\text{g}}^{3}$ approaches 00, from the Collapsed phase where $\rho$ is finite, see the green line Figure 3A. This transition is continuous but sharp in the large $N_{\text{p}}$ limit  [14, 20].
@@
Here $t_{\text{b}}$ corresponds to the reduced temperature, a measure of interaction strength between bulk molecules, $\mu_{\text{b}}$ is the chemical potential, and $u$ captures higher order interactions and is required for stability. For $t_{\text{b}}<0$, coexistence occurs at $\mu_{\text{b}}=0$ with energy minima located at $\phi=\pm\sqrt{6t_{b}/u}$. For $t_{\text{b}}\geq 0$ there is one stable phase set by $\mu_{\text{b}}$. These two phases meet in a critical point at $t_{\text{b}}=0$.
In a standard mean field approximation without a polymer, the order parameter is assumed to take the value $\phi_{\infty}$ which globally minimizes the free energy density $f_{\text{bulk}}$. In the presence of a long polymer, we further assume that a sphere of volume $R_{\text{g}}^{3}$ has a polymer density $\rho$ and a potentially different value of the order parameter $\phi$. Here we assume that both order parameters, $R_{\text{g}}$ (or equivalently $\rho$) and $\phi$ together minimize the free energy of the coupled system.
We describe interactions between the bulk and the long polymer via an interaction energy density $f_{\text{int}}$ and corresponding interaction energy $F_{\text{int}}=f_{\text{int}}R_{\text{g}}^{3}$:
@@
These interactions can be interpreted as the binding energy between the bulk and the long-polymer. To compactly write the coupled free energy of the polymer and its order parameter environment we define $\Delta f_{\text{bulk}}=f_{\text{bulk}}(\phi)-f_{\text{bulk}}(\phi_{\infty})$. The free energy of the coupled system can then be written as:
@@
A super-critical bulk system modifies the polymer collapse transition — To see how the bulk modifies the thermodynamics of the long polymer, we examine the behavior of a weakly interacting bulk by approximating the free energy as quadratic, $\Delta f_{\text{bulk}}=\frac{t_{\text{b}}}{2}\left(\phi-\phi_{\infty}\right)^{2}$. In this quadratic approximation the bulk can be minimized over $\phi$ analytically, leaving just one additional $R_{\text{g}}$ dependant contribution to the effective free energy of an isolated polymer, so that $F^{\text{eff}}_{\text{poly}}(R_{\text{g}})=F^{0}_{\text{poly}}(R_{\text{g}})-\frac{h^{2}}{t_{\text{b}}}N_{\text{p}}\rho$, and a constant term $-h\phi_{\infty}N_{\text{p}}$. This additional term has exactly the same form as the monomer-monomer coupling, defining a new $v_{\text{eff}}\equiv v_{\text{p}}-\frac{h^{2}}{t_{\text{b}}}$. Since there are no other qualitative changes to the free energy, the effect of the bulk in this regime is precisely to shift the parameter regime of the collapse transition, see purple line in Figure 3A. The value of the bulk order parameter tracks with the density of the polymer and sees a sharp, continuous increase through the collapse transition, see Figure 3B, purple line. This effective interaction becomes very strong near the critical point of the bulk, where $t_{\text{b}}$ is near zero.
Figure 3: Bulk phase-transitions modify the polymer collapse transition in a coupled mean field theory: (A) Free energy density of the bulk, $\Delta f_{\text{bulk}}$, in super-critical (purple) and sub-critical (blue) regimes. We distinguish these regimes by the presence of one minimum or two well-separated minima. (B) Polymer density $\rho$ as a function of monomer-monomer interaction strength $v_{\text{p}}$. An isolated polymers sees a continuous but sharp collapse transition at $v_{\text{p}}=0$ (green line), but in coupled systems collapse occurs at repulsive $v_{\text{p}}>0$ in continuous (purple, bulk has one minima) or discontinuous (blue, bulk has two minima) transitions. (C) Bulk density difference, $\phi-\phi_{\infty}$, as a function of bulk chemical potential $\mu_{\text{b}}$. In coupled systems $\phi-\phi_{\infty}$ increases with $\mu_{\text{b}}$ in continuous (purple) and discontinuous (blue) transitions. In an uncoupled system bulk condensation occurs at $\mu_{\text{b}}=0$ (not visible on this plot). (D) (left) Quadratic approximation to bulk free energy density in the subcritical regime. Orange corresponds to Dilute phase, purple to Dense phase, and dashed blue is the full free energy. (top) Free energy landscape for a system in the Extended, Dilute phase. Local minima corresponding to Extended (orange) and Collapsed (purple) phases are marked. The absolute minima of $\phi-\phi_{\infty}$ (dashed, blue) interpolates between the quadratic approximations (solid lines, purple and orange). (bottom) Free energy as a function of polymer density $\rho$. The global minimum of this system is at the low-density, low-$\phi$ minima (orange). E) Energy landscape (above) and free energy as a function of polymer density (below) for a system in the Collapsed, Prewet phase. The global minimum of this system occurs at the high density, high $\phi$ minima.

### v2

# Protein-DNA Co-condensation is Prewetting to a Collapsed Polymer

@@ Abstract
The three-dimensional organization of chromatin is thought to play an important role in controlling gene expression. Specificity in expression is achieved through the interaction of transcription factors and other nuclear proteins with particular sequences of DNA. At unphysiological concentrations many of these nuclear proteins can phase-separate in the absence of DNA. In-vivo the thermodynamic forces driving these phases lead the chromosome to co-condense with nuclear proteins. However it is unclear how DNA, itself a long polymer subject to configurational transitions, interacts with three-dimensional protein phases. Here we show that a long compressible polymer can be coupled to interacting protein mixtures, leading to a generalized prewetting transition where polymer collapse is coincident with a locally stabilized liquid droplet. We use lattice Monte-Carlo simulations and a mean-field theory to show that these phases can be stable even in regimes where both polymer collapse and coexisting liquid phases are unstable in isolation, and that these new transitions can be either abrupt or continuous. For polymers with internal linear structure we further show that changes in the concentration of bulk components can lead to changes in three-dimensional polymer structure. In the nucleus there are many distinct proteins that interact with many different regions of chromatin, potentially giving rise to many different Prewet phases. The simple systems we consider here highlight chromatin’s role as a lower-dimensional surface whose interactions with proteins are required for these novel phases.
Significance Statement — Proteins and DNA in the nucleus display rich spatial organization, but the forces which drive it are not well understood. Here we show that proteins prone to forming condensed liquid droplets can drive configurational phase transitions of long polymers, like DNA, even when too dilute to phase separate on their own. Indeed, many transcription factors (TFs) will condense into liquid phases in the absence of DNA when enriched to much higher concentrations than in the nucleus. With DNA, and at much lower TF concentrations, we expect these proteins to undergo generalized prewetting transitions, leading to abrupt changes in the three dimensional organization of chromatin. We argue that these phase transitions play an important role in organizing and regulating chromatin.
@@
In the nucleus of eukaryotes chromosomes display three dimensional structure at many distinct scales [1]. In some cases the co-localization of enhancers, promoters, and otherwise distal genetic elements has been shown to be a key step in initiating transcription [2, 3, 4]. While some of this structure is driven by enzymatic activity from loop extrusion factors (LEFS) and polymerases  [5, 6], some is likely to be thermodynamic in origin, driven by energetic interactions between long polymers and proteins, small RNAs and other macromolecules [7, 8, 9]. Many nuclear components have a tendency to phase separate into coexisting liquid phases when isolated at high concentrations, even in the absence of DNA [10, 7, 11]. These protein phases likely co-condense with chromatin to organize the genome  [12, 13, 2, 14, 15].
@@
But there are other thermodynamic drivers of spatial organization available to long polymers like chromosomes. In particular, Collapse transitions separate a regime where typical configurations are extended from one where polymers are condensed and space filling [16]. These transitions can be accessed by solvent quality; in good solvents polymers typically have a weak contact repulsion and exist in an Extended configurational phase. In poor solvents, polymers prefer to maximize self-contacts, leading to Collapsed phases. In the nucleus more complex interactions could in principle drive more interesting configurational phases.
@@
We recently argued that a range of structures at the plasma membrane are Prewet: surface phases formed by a combination of membrane mediated forces and interactions between cytoplasmic proteins [17]. In classical prewetting, a thin film resembling a three dimensional phase can adhere to a surface outside of a thermodynamic regime where that bulk phase is stable [18, 19]. But in the biological case, the solid surface is replaced by the fluid plasma membrane, itself found near a 2D liquid-liquid demixing critical point [20]. In this system the prewetting transition of the bulk merges with the de-mixing transition at the surface, allowing for phase transitions that would not occur in the membrane or bulk alone.
@@
Here we explore a model in which bulk proteins prone to phase separation play an analogous role in modulating the properties of a polymer collapse transition. In this generalized prewetting transition there is a phase in which bulk components are at near condensed density, and in which the polymer is Collapsed, at parameters in which the polymer would remain Extended in the absence of bulk, and the bulk would remain well-mixed in the absence of a polymer. We investigate the phase diagram of this generalized prewetting transition with both a lattice model and a mean field theory.
@@
Model: We model a single self-avoiding polymer on a $3$-D cubic lattice of linear dimension $L=64$. A spin variable at lattice site $i$, $s^{p}_{i}=1$ for sites which the polymers passes through and is 00 otherwise. The energy of configurations of this system is given by the Hamiltonian $H_{\text{poly}}$:
@@
Here both sums are over lattice sites, and $<i,j>$ represents nearest neighbors. $\mu_{p}$ is a monomer chemical potential which multiplies $N_{\text{p}}=\sum_{i}s_{i}^{\text{p}}$, the total number of monomers in the polymer. $v_{\text{p}}$ is a monomer-monomer interaction energy, and the subtraction in the first term avoids counting contacts between sequential monomers. All simulations take place on periodic lattices, and are constrained to polymer configurations with initial position $(L/2,L/2,0)$ and final position $(L/2,L/2,L)$. We also enforce that the polymer have length shorter than $N_{\text{max}}$, typically $1500$ monomer units, so that the polymer obeys $L\leq N_{\text{p}}<N_{\text{max}}$. In most of our simulations $N_{\text{p}}\approx N_{\text{max}}$, so that we sample an effectively fixed length distribution, but allow small fluctuations for numerical convenience.
We couple this single long polymer (yellow in Fig.1A) to a solution of two types of shorter polymers which we term bulk molecules (red and blue in Fig.1A). Bulk molecules are held at fixed chemical potential $\mu_{\text{b}}$ and fixed length $N_{\text{b}}$. Following our previous work [17], we model the bulk as a mixture of two types of molecules prone to condensing together. Molecules of the same type are self-avoiding while molecules of different types may occupy the same site, and have an on-site attractive interaction with energy $J_{\text{bulk}}$. Their occupancies at site $i$ are defined analogously to the long polymer by two additional spin variables $s^{1}_{i}$ and $s^{2}_{i}$. All molecules interact with a weak nearest-neighbor energy $J_{\text{nn}}$. The bulk Hamiltonian thus reads:
$\frac{H_{\text{bulk}}}{k_{\text{B}}T}=$ $-\frac{\mu_{\text{b}}}{N_{\text{b}}}\sum_{i}\left(s^{1}_{i}+s^{2}_{i}\right)-J_{\text{bulk}}\sum_{i}s^{1}_{i}s^{2}_{i}$
$-J_{\text{nn}}\sum_{<i,j>}\left(s^{1}_{i}s^{1}_{j}+s^{1}_{i}s^{2}_{j}+s^{2}_{i}s^{2}_{j}\right)$ (2)
@@
We illustrate this model in Fig.1A.
@@
We use a Monte-Carlo procedure to sample equilibrium configurations of this system (see Methods). We sample the long polymer through three elementary moves that add, delete, or ‘kink’ a given bond. We exchange bulk molecules with a particle reservoir and update their configuration with moves that conserve length.
Figure 2: The bulk system shifts the Polymer Collapse phase boundary — (A-D) Configurations of coupled phases: Outside of the bulk coexistence region there are two phases that correspond to phases of an isolated polymer: (A) Extended and (B) Collapsed. The Collapsed phase coincides with condensation, or Prewetting, of the bulk fluid. The Extended phase has corresponding Dilute bulk phase. Within the bulk coexistence region the polymer may prefer to enter the droplet, confining itself (Wet, C) or avoiding it (Dry, D). In C and D we hold bulk particle number fixed to observe bulk coexistence. (E) Schematic bulk phase diagram: Bulk coexistence of a dense and dilute phase occurs in the yellow shaded region. Generalized Prewet phases that coincide with polymer collapse occur outside of the bulk coexistence region (light blue region). These can be accessed by increasing bulk density through a first order transition (solid blue line) or through a continuous transition (dashed blue line). These lines meet at a tricritical point (blue x). Colored points correspond to above configurations. (F) Schematic phase diagram of monomer-monomer interactions $v_{\text{p}}$ and bulk chemical potential $\mu_{\text{b}}$. The bulk modifies the phase boundary of the polymer, allowing it to collapse in regions where it otherwise would not; this Collapse transition coincides with condensation of the bulk on the surface of the polymer.
@@
Thermodynamics of an Isolated Polymer — In the absence of the bulk system, the long polymer can access several thermodynamic phases. These phases correspond to distributions over the spatial configurations of the long polymer, $\left\{\vec{r}_{i}\right\}$, where $\vec{r}_{i}$ is the position of the $i^{\text{th}}$ monomer. For idealized polymers in regions where boundary conditions can be neglected, these phases can be characterized by scaling relationships between $N_{\text{p}}$ and the radius of gyration $R_{\text{g}}$, defined by $R_{\text{g}}^{2}=1/N_{\text{p}}\sum_{i}^{N_{\text{p}}}\left(\vec{r}_{i}-\vec{r}_{\text{cm}}\right)^{2}$ with $\vec{r}_{cm}$ the center of mass of the chain. We show sample configurations of these phases in Fig.1B
@@
When $\mu_{p}$ is negative and $v_{\text{p}}$ is positive or small and negative, the polymer is in the Short phase. Here $N_{\text{p}}$ is small, $\approx L$. In our simulations we typically set $\mu_{\text{p}}=0$, and are not concerned with this Short phase.
@@
When $v_{\text{p}}$ is large and attractive, the polymer is in the Collapsed phase. Here $N_{\text{p}}$ fluctuates about $N_{\text{max}}$, and typical configurations are space-filling, with a constant density as $N_{\text{max}}\rightarrow\infty$.
@@
When $v_{\text{p}}$ is large and repulsive, but $\mu_{p}$ is positive, the polymer is in the Extended phase. Here the length fluctuates about $N_{\text{max}}$, but the polymer configuration is Extended and the average density of the polymer approaches 00 as $N_{\text{max}}$ approaches infinity.
The Extended and Collapsed phase are separated by a line of continuous transitions at $v_{\text{p}}=0$. We show a typical Extended configuration in Fig.1B.
@@
Bulk Molecules Phase Separate and Modify the Polymer Phase Diagram — Our bulk system has phases and phase-transitions independent of its interactions with the long polymer. The bulk phases are dilute and dense corresponding to gas and liquid-like states. At certain chemical potentials, or densities, these bulk phases will coexist and the system phase-separates, drawn as the yellow region in Fig.2E, with typical configurations in each phase shown in Fig.1C.
@@
The polymer configurations in the Wet phase depend on $J_{\text{int}}$ and $J_{\text{bulk}}$, which together determine solvent quality of the dense phase. A polymer that prefers to be Extended in an infinite bulk dense phase can be confined by a finite-size droplet separate from the collapse transition.
@@
Outside of the bulk coexistence region we find another phase where the polymer is Collapsed and also enriched with a dense-like phase of bulk molecules, shown in Fig.2B. This is reminiscent of prewetting where a thin bulk domain forms which is adhered to a particular surface phase [17, 19]; we call phases where polymer collapse coincides with polymer-only bulk de-mixing Generalized Prewet - Collapsed. Consistent with this view, the polymer phase-diagram effectively shifts and the collapse transition occurs at repulsive $v_{\text{p}}$. Polymer collapse and bulk demixing couple because this minimizes the volume of an otherwise unfavorable bulk phase, while maintaining a high density of bulk-bulk and bulk-polymer interactions. Increasing $\mu_{\text{b}}$, we transition discontinuously between the three phases shown in Fig.2A-C, which we draw schematically in the phase diagrams sketched in Fig.2E,F.
@@
Here we develop a mean-field theory that couples a long, collapsible polymer to a bulk fluid with a propensity to phase-separate.
@@
The first two terms labeled ‘Non-Interacting polymer’ are entropic contributions to the free-energy from gaussian statistics, and can be derived by replacing the radius of gyration with the end-to-end distance, with $a$ the linear dimension of a single monomer. A free-energy written in terms of $R_{\text{g}}$ has additional entropic costs due to confinement – but these, along with higher order interactions, are never relevant so we neglect them here (see Supplement). The final two terms capture mean-field interactions between monomers [21]. The parameter $v_{\text{p}}$ is the difference between monomer-monomer and monomer-solvent interactions. The point $v_{\text{p}}=0$ separates an Extended polymer phase, where the monomer density $\rho=Na^{3}/R_{\text{g}}^{3}$ approaches 00, from the Collapsed phase where $\rho$ is finite, see the green line Fig.3B. This transition is continuous but sharp in the large $N_{\text{p}}$ limit  [16, 22].
Figure 3: Bulk phase-transitions modify the polymer collapse transition in a coupled mean field theory: (A) Free energy density of the bulk, $\Delta f_{\text{bulk}}$, in super-critical (purple) and sub-critical (blue) regimes. We distinguish these regimes by the presence of one minimum or two well-separated minima. (B) Polymer density $\rho$ as a function of monomer-monomer interaction strength $v_{\text{p}}$. An isolated polymers sees a continuous but sharp collapse transition at $v_{\text{p}}=0$ (green line), but in coupled systems collapse occurs at repulsive $v_{\text{p}}>0$ in continuous (purple, bulk has one minima) or discontinuous (blue, bulk has two minima) transitions. (C) Bulk density difference, $\phi-\phi_{\infty}$, as a function of bulk chemical potential $\mu_{\text{b}}$. In coupled systems $\phi-\phi_{\infty}$ increases with $\mu_{\text{b}}$ in continuous (purple) and discontinuous (blue) transitions. In an uncoupled system bulk condensation occurs at $\mu_{\text{b}}=0$ (not visible on this plot). (D) (left) Quadratic approximation to bulk free energy density in the subcritical regime. Orange corresponds to Dilute phase, purple to Dense phase, and dashed blue is the full free energy. (top) Free energy landscape for a system in the Extended, Dilute phase, $v_{\text{p}}=2$. Local minima corresponding to Extended (orange) and Collapsed (purple) phases are marked. The absolute minima of $\phi-\phi_{\infty}$ (dashed, blue) interpolates between the quadratic approximations (solid lines, purple and orange). (bottom) Free energy as a function of polymer density $\rho$. The global minimum of this system is at the low-density, low-$\phi$ minima (orange). (E) Energy landscape (above) and free energy as a function of polymer density (below) for a system in the Collapsed, Prewet phase, $v_{\text{p}}=0.1$. The global minimum of this system occurs at the high density, high $\phi$ minima.
@@
Here $t_{\text{b}}$ corresponds to the reduced temperature, a measure of interaction strength between bulk molecules, $\mu_{\text{b}}$ is the chemical potential, and $u$ captures higher order interactions and is required for stability.
In a standard mean field approximation without a polymer, the order parameter is assumed to take the value $\phi_{\infty}$ which globally minimizes the free energy density $f_{\text{bulk}}$. In the presence of a long polymer, we further assume that a sphere of volume $R_{\text{g}}^{3}$ has a polymer density $\rho=\frac{N_{\text{p}}a^{3}}{R_{\text{g}}^{3}}$ and a potentially different value of the order parameter $\phi$. Here we assume that both order parameters, $R_{\text{g}}$ (or equivalently $\rho$) and $\phi$ together minimize the free energy of the coupled system.
We describe interactions between the bulk and the long polymer via an energy density $f_{\text{int}}$ and corresponding interaction energy $F_{\text{int}}=f_{\text{int}}R_{\text{g}}^{3}$:
@@
The parameter $h$ is analogous to $J_{\text{int}}$, and can be interpreted as a per monomer binding interaction between bulk and long polymer. To compactly write the coupled free energy of the polymer and its order parameter environment we define $\Delta f_{\text{bulk}}=f_{\text{bulk}}(\phi)-f_{\text{bulk}}(\phi_{\infty})$ and write the free energy of the coupled system as:
@@
A super-critical bulk system modifies the polymer collapse transition — To see how the bulk modifies the thermodynamics of the long polymer, we examine the behavior of a weakly interacting bulk by approximating the free energy as quadratic, $\Delta f_{\text{bulk}}=\frac{t_{\text{b}}}{2}\left(\phi-\phi_{\infty}\right)^{2}$. In this approximation $\phi$ can be minimized analytically, leaving just one additional $R_{\text{g}}$ dependent contribution to the effective free energy of an isolated polymer, so that $F^{\text{eff}}_{\text{poly}}(R_{\text{g}})=F^{0}_{\text{poly}}(R_{\text{g}})-\frac{h^{2}}{t_{\text{b}}}N_{\text{p}}\rho$, and a constant term $-h\phi_{\infty}N_{\text{p}}$. This additional term has exactly the same form as the monomer-monomer coupling, defining a new $v_{\text{eff}}\equiv v_{\text{p}}-\frac{h^{2}}{t_{\text{b}}}$. Thus, the effect of the bulk in this regime is precisely to shift the parameter regime of the collapse transition, see purple line in Fig.3A. The value of the bulk order parameter tracks with the density of the polymer and sees a sharp, continuous increase through the collapse transition, see Fig.3B, purple line.